OpenStudy (anonymous):

solve the problem m/1-m+1+m/m=13/6

OpenStudy (anonymous):

(m)/(1)-m+1+(m)/(m)=(13)/(6) Find the LCD (least common denominator) of (m)/(1)-m+1+(m)/(m)+(13)/(6). Least common denominator: 6m Multiply each term in the equation by 6m in order to remove all the denominators from the equation. (m)/(1)*6m-m*6m+1*6m+(m)/(m)*6m=(13)/(6)*6m Simplify the left-hand side of the equation by canceling the common terms. 12m=(13)/(6)*6m Simplify the right-hand side of the equation by simplifying each term. 12m=13m Since 13m contains the variable to solve for, move it to the left-hand side of the equation by subtracting 13m from both sides. 12m-13m=0 Since 12m and -13m are like terms, add -13m to 12m to get -m. -m=0 Multiply each term in the equation by -1. -m*-1=0*-1 Multiply -m by -1 to get m. m=0*-1 Multiply 0 by -1 to get 0. m=0 Verify each of the first set of solutions by substituting them into the original equation and solving. In this case, none of the solutions are valid. No Solution